OpenStudy (anonymous):

Solve the equation 1/x-4 = x/ x^2-16

OpenStudy (anonymous):

Can you write the equation explicitly ..

OpenStudy (anonymous):

What do you mean?

OpenStudy (anonymous):

Is it\[\frac{ 1 }{ x-4 } = \frac{ x }{ x^2 - 16}\]

OpenStudy (anonymous):

This is the problem

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

Use x^2-16 = (x-4)*(x+4)

OpenStudy (anonymous):

Would I then times x+4 to both sides?

OpenStudy (anonymous):

Cross multiply and factorize.

OpenStudy (anonymous):

Actually take (x-4) as common.

OpenStudy (anonymous):

When I cross multiply I get (x-4)(x+4)=x^2-4 Is that correct?

OpenStudy (anonymous):

(x-4)(x+4)=x^2-4x

OpenStudy (anonymous):

Make right hand side 0

OpenStudy (anonymous):

Wait I just got confused

OpenStudy (triciaal):

@ShailKumar (x-4)(x+4) = x^2-16

OpenStudy (anonymous):

Manipulation of the equation gives: (x-4)(x+4)=x^2-4x